OpenStudy (anonymous):

Hypothesis Testing - A claim is made that the mean of a set of data will be 4.7. In a test of the claim, the mean after 49 trials is 5, with a standard deviation of 0.5. Which of the following numbers is the z-value of the given data? Round to the nearest hundredth.

OpenStudy (amistre64):

define Ho and Ha in order to run the claim setup

OpenStudy (amistre64):

or if you got a ti83 ... run a z interval

OpenStudy (amistre64):

z-test that is

OpenStudy (amistre64):

\[z=\frac{\bar x-[claim]}{\sigma \sqrt n}\]

OpenStudy (anonymous):

\[z = \frac{ x - 2 }{ \frac{ o }{ \sqrt{n} }}\] all i know is this so far

OpenStudy (amistre64):

5-4.7 , not x-2

OpenStudy (anonymous):

i meant the u sign

OpenStudy (anonymous):

how do i know what i plug in where ?

OpenStudy (amistre64):

the claim is 4.7 so we are using that as a mean to test from